Comparing the tax on $350,000 of taxable income for each tax filing status, the filing status that produces the greatest amount of tax is married filing seperatly. True of False
Answer #1

Answer: False

[Based on the US tax rates for the year 2018]

A single filer:

Tax amount = 45,689.50 + (350,000 – 200,000 =) 150,000 × 35%

                        = 45,689.50 + 52,500

                        = 98,189.50

Married – filing jointly:

Tax amount = 64,179 + (350,000 – 315,000 =) 35,000 × 32%

                        = 64,179 + 11,200

                        = 75,379

Married – filing separately:

This is same as “a single filer”.

Tax amount = 45,689.50 + (350,000 – 200,000 =) 150,000 × 35%

                        = 45,689.50 + 52,500

                        = 98,189.50

Head of household:

Tax amount = 44,298 + (350,000 – 200,000 =) 150,000 × 35%

                        = 44,298 + 52,500

                        = 96,798

Therefore, the tax amount of “married – filing separately” doesn’t have the highest value, since “a single filer” is also paying the same tax amount.